OpenStudy (anonymous):

What is the solution to the system of equations that contains –3x + y = 3 and 2x – y = –1? A. no solution B. (–1, 3) C. infinite number of solutions D. (–2, –3

OpenStudy (anonymous):

add them together –3x + y = 3 2x – y = –1 ---------- -x = 2

OpenStudy (anonymous):

thanks but just to make sure the answer is b right

OpenStudy (anonymous):

how do you get conclusion it's b ?

OpenStudy (anonymous):

well I see –3x + y = 3 2x – y = –1 and the last numbers for both those equations are the same #'s as B

OpenStudy (anonymous):

I believe the ( , ) states values that x and y must take in our case, we found x = -2 and there was no constraint for y there are probably infinitely many solutions

OpenStudy (anonymous):

you can look at it this way -> no matter what we choose for y: the second equation always cancels the first one

OpenStudy (anonymous):

oh :) ok thanks

OpenStudy (anonymous):

you're welcome ;)

OpenStudy (anonymous):

The answer ended up being D

OpenStudy (anonymous):

thanks for the feedback

OpenStudy (anonymous):

haha no problem

OpenStudy (anonymous):

–3(-2) + y = 3 6 + y = 3 y = 3-6 y=-3 2(-2) – y = –1 -4 -y = -1 -4 = y -1 -3 = y

OpenStudy (anonymous):

I believe there are infinitely many solutions if you get 0=0 in case you run into a problem like that in the future

OpenStudy (anonymous):

-3x + y = 3 2x - y = -1 ----------- -x = 2 x = -2 now sub -2 in for x in either of the equations 2x - y = -1 2(-2) - y = -1 -4 - y = -1 -y = -1 + 4 -y = 3 y = -3 answer is D.